56. help--
A bus leave a station at 1 p.m. , traveling west an average rate of 44 mi/h. One hour later a second bus leaves the same station,traveling each at a rate of 58 mi/h.At what time will the two buses be 274 mi apart?

Answer by Nate(3500)

First bus:
y = 44x + 44
Second bus:
q = 58x
Combine:
q - y = 274
58x - (44x + 44) = 274
58x - 44x - 44 = 274
14x = 318
x = 159/7 or about 22.7 hours
22 hours and 42 minutes; means 1 hour and 18 minutes from a whole day
11:42 a.m.